# A Comprehensive Survey on Vector Database: Storage and Retrieval Technique, Challenge

### v2

@@ III-B4 Quantization-Based Approach
The O-PQ algorithm also updates the codebooks and codes for each subvector using an online learning technique. The algorithm uses two parameters: $\alpha$, which controls the learning rate, and $\beta$, which controls the forgetting rate. The algorithm updates the codebooks and codes as follows:
For each new point $x$, assign it to its nearest centroid in each subvector using PQ.
For each subvector $x_{i}$, update its centroid $c_{q_{i}\left(x_{i}\right)}$ as:
@@
The advantage of O-PQ is that it can deal with changing data distributions and new data points, as it can update the codebooks and the codes in real time.
O-PQ search performance depends on roughly the same factors as OPQ, and also faces similar challenges as OPQ.
Scalable Nearest Neighbor. It is a technique for efficient vector similarity search at scale[96, 101]. ScaNN optimizes Maximum Inner Product Search (MIPS) through search space pruning and quantization. Traditional MIPS schemes aim to minimize the average distance between each vector $x$ and its centroids $\tilde{x}$, that is, to minimize quantization distortions.
The formula for typically measure the quantization distortion is:
@@
where $N$ is the total number of vectors, $x_{i}$ is original vector, $\tilde{x}$ is quantized centroid, and $\|\cdot\|_{2}$ denotes Euclidean norm.
@@
ScaNN taking into account the distribution characteristics of the data in different directions, ellipsoidal or other shaped regions are used instead of spherical regions around the centroids to better fit the local structure of the data. Building on this perspective, the anisotropic loss function can further enhance the adaptability of vector quantization to data anisotropy. By explicitly separating quantization errors into parallel and orthogonal components, the anisotropic loss function assigns distinct scaling parameters $h_{i}^{\|}$ and $h_{i}^{\perp}$ to these components, respectively. This allows for more fine-grained control over the quantization process, ensuring that the errors are distributed in alignment with the data’s geometric characteristics.
@@
By integrating this anisotropic loss framework, the quantization process moves beyond spherical symmetry and better accommodates ellipsoidal or irregularly shaped distributions in the data. This alignment with ScaNN’s approach to using ellipsoidal regions around centroids enhances both the representation accuracy and the retrieval efficiency, especially in cases where the data exhibits significant directional variance.
Moreover, extending this concept to product quantization allows the construction of multiple subspace-specific dictionaries, each tuned to the anisotropic characteristics of the corresponding subspace. This not only retains the efficiency of ScaNN’s design but also adds a layer of flexibility for handling complex, high-dimensional data distributions. The performance of ScaNN depends on several factors, such as the anisotropy of the data distribution, the choice of quantization methods like vector or product quantization, the size and quality of the codebooks, and the efficiency of the partitioning and scoring processes. These factors impact the trade-off between retrieval accuracy and computational speed, as well as the scalability of the algorithm for massive datasets. Additional challenges and potential extensions include handling highly irregular data distributions, selecting optimal scaling parameters for the score-aware loss function, and integrating ScaNN with complementary techniques like hierarchical search structures or advanced compression methods.
@@
Inverted File Product quantization. It is a widely used technique for approximate nearest neighbor (ANN) search in high-dimensional vector spaces[40, 97]. This algorithm is a combination of the Inverted File Indexing (IVF) and Product Quantization (PQ) algorithms. IVF_PQ first uses the IVF algorithm to divide or partition the data into clusters and uses the parameter $nprobe$ to control the number of clusters. The higher the $nprobe$, the better the search results, but it also increases the time required. It then identifies the top-N clusters closest to the query vector and performs the search within these N clusters using the Product Quantization (PQ) algorithm.
The IVF_PQ algorithm naturally results in two different approaches when using the PQ algorithm: the first involves performing K-means clustering with the IVF algorithm, followed by applying a local PQ algorithm for dimensionality reduction within each cluster; the second also starts with the IVF algorithm to divide all data points into several clusters but applies a globally unified PQ algorithm for dimensionality reduction within each cluster.In addition, there is another implementation scheme for the IVFPQ algorithm in FAISS (Facebook AI Similarity Search) [^1] https://ai.meta.com/tools/faiss/. First, all data points are clustered using the IVF algorithm. Then, for all data points within each cluster, the difference between each point and its cluster center (referred to as the ”””residual”) is calculated. Mathematically, the residual represents the offset of a data point relative to its cluster center. This operation is equivalent to shifting all cluster centers to the origin, causing all points to focus around the origin. Afterward, the PQ algorithm is applied to the residuals. Compared to the previous two approaches, the key difference is that PQ is applied to the residuals rather than the original vectors. The advantage of this scheme is that, as the data points become more tightly clustered, the average size of each cluster region is smaller, leading to reduced approximation errors during distance computations.
IV Vector Database Comparison
@@
IV-1 The Comparison of Features and Characteristics of Vector Databases
Table I: FEATURES OF VECTOR DATABASES
@@ IV-1 The Comparison of Features and Characteristics of Vector Databases
The characteristics of VDBs directly affect their performance in practical applications. Therefore, gaining a deep understanding of these databases’ features is essential for selecting the most suitable one. As shown in Table I, we compare several popular VDBs, focusing on their differences in indexing methods, query types, distance functions, scalability, maximum dimension, and support for data management features such as replication, sharding, and partitioning.
It can be observed from the table I that all VDBs support NNS and ANNS. However, the implementation strategies and optimizations for these searches vary significantly across databases, depending on their underlying indexing methods and architectural designs. For example, the indexing methods and distance functions are not exactly the same across databases, but there are commonalities. For instance, all databases except Pinecone support graph-based methods, which indicates that graph-based methods are widely adopted for their ability to handle complex relationships and data structures. Additionally, the majority of databases also support three distance functions: inner product, cosine similarity, and Euclidean distance. For details on the indexing methods and distance functions supported by different databases, see the table II and table III below.
Table II: Overview of Supported Distance Functions in Vector Databases
@@
Table III: Overview of Supported Indexing Methods in Vector Databases
@@
$SPARSE\_Inverted\_Index$$\times$ $\times$ $\times$ $\times$ $\times$ $\times$ N/A
$SPARSE\_WAND$$\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ N/A
@@
The last column of the table I provides statistics on the maximum vector dimensions supported by each database. It can be observed that, with the exception of ElasticCloud and Pinecone, for which no relevant information was available, most of the listed VDBs support a total vector dimension in the range of tens of thousands, with the maximum supported dimensions ranging from 16,000 to 65,535. It should be noted that QdrantCloud has a default support for up to 65,535 dimensions, though this can be configured to support higher dimensions.
IV-2 The Comparison of Loading Capacity and Search Performance of Vector Database
@@
VectorDBBench provides a comprehensive performance analysis by evaluating VDBs based on metrics such as Queries Per Second (QPS), recall rate, latency(the time required for each query from submission to system response), load duration, and maximum load count(The maximum number of vectors a database can successfully insert or store in a single loading operation.). Its testing methodology employs a relative scoring mechanism to ensure fair comparisons. For QPS, the highest observed value among all tested databases serves as the reference baseline; for latency, the lowest observed value among all tested databases is used as the baseline, with an additional 10ms adjustment to avoid distortions when latency is very low. For systems that fail or encounter timeouts in a specific test case, their scores are penalized by assigning a value proportionally worse than the lowest-performing result, using a factor of two. For example, in the case of QPS, the score is reduced to half of the minimum observed value, while for latency, it is increased to twice the maximum observed value. The formulas for calculating QPS and latency metrics for VDB $x$ are as follows:
@@
where $origin\_QPS_{x}$ and $origin\_Latency_{x}$ represent the original QPS value and original latency value, respectively, measured for database $x$ during the test. $base\_QPS$ and $base\_Latency$ is the reference baseline.
Table IV: VECTOR DATABASE EVALUATION TEST CASES
@@ 3 Search Performance Google C43 500K 1536 N/A IBT, R, L, MQPS
Specifically, as shown in table IV, the VDB evaluation consists of a series of test cases designed to assess capacity, search performance, and filtering search performance. Capacity cases (Cases 1 and 2) measure the database’s ability to handle large datasets, focusing on the number of inserted vectors using SIFT and GIST datasets. Search performance cases (Cases 3) evaluate index building time, recall, latency, and maximum QPS using Google C4 dataset dataset.
Figure 7: Performance Test Results for Vector Databases
@@
The VDB versions involved in the performance tests are as follows: Milvus-2c8g-hnsw-v2.2.12 (hereafter referred to as Milvus), Pinecone-p1.x1 (hereafter referred to as Pinecone), WeaviateCloud-standard (hereafter referred to as Weaviate Cloud), ZillizCloud-2cu-cap-v2023.6 (hereafter referred to as ZillizCloud), QdrantCloud-2c8g-1node (hereafter referred to as QdrantCloud), PgVector-2c8g (hereafter referred to as PgVector), and ElasticCloud-upTo2.5c8g (hereafter referred to as ElasticCloud). To ensure minimal differences in hardware performance across the tested databases, a configuration of 2 CPUs and 8GB of memory was specifically selected. For VDBs that do not meet this hardware requirement, similar configurations were chosen as closely as possible. The specific test results are shown in the figure7. The overall ranking in the figure 7 is calculated by averaging the rankings of each sub-test item, with the final overall ranking determined in ascending order of the average values. According to the overall ranking, Milvus ranks first with its comprehensive performance, achieving a QPS of 380, latency of 12.4 milliseconds, and outstanding performance in load capacity. In contrast, ElasticCloud ranks last, with a QPS of 11.29 and latency as high as 361 milliseconds. In terms of recall rate, all databases perform close to 1.0, indicating little difference in query accuracy. ZillizCloud demonstrates the best latency performance, at only 6 milliseconds, but its load capacity is relatively low. It is worth noting that ZillizCloud, ElasticCloud, and PgVector were assigned penalty values (P) due to timeouts in the load capacity tests, which may have been caused by network issues and should not be taken as a definitive measure of their actual performance. Overall, no single database ranked in the top three across all tests, indicating that different databases may have their own strengths. Additionally, the overall rank is simply a straightforward average of the rankings in each test item, without applying weighted averages based on specific tasks, so it should be regarded as a preliminary reference.
V Challenges
V-A Index Construction and Searching of High-Dimensional Vectors
@@
VDBs require efficient indexing and searching of billions of vectors in hundred or thousand dimensions, which poses a huge computational and storage challenge. Traditional indexing methods, such as B-trees or hash tables, are not suitable for high-dimensional vectors because they suffer from dimensionality catastrophe. Therefore, VDBs need to use specialized techniques such as ANN search, hashing, quantization, or graph-based search to reduce complexity and improve the accuracy of vector similarity search.
V-B Support for Heterogeneous Vector Data Types
@@
V-C Distributed Parallel Processing Support
@@
V-D Integration with Mainstream Machine Learning Frameworks
VDBs need to be integrated with popular machine learning frameworks such as TensorFlow, PyTorch, Scikit-learn, etc., which are used to generate and use vector embeddings. As a result, VDBs need to provide easy-to-use application programming interfaces (API) and encapsulated connectors to seamlessly interact with these frameworks and support a variety of data formats and models.
V-E Integration with Emerging Application Scenarios
@@
Currently, many emerging application scenarios remain underexplored. For instance, the incremental k-NN search[102] adopted by recommendation and e-commerce platforms faces significant challenges due to the vast imbalance between the volume of vector data processed and the data displayed to users. This method cannot be effectively supported by most VDBs[20]. Furthermore, with the latest advancements in sparse vector technology, integrating these technologies into VDBs to enable hybrid retrieval (combining keyword and vector retrieval methods) is increasingly regarded as a best practice. Such hybrid systems must manage large datasets while enhancing computational efficiency and maintaining retrieval quality.
V-F Data Security and Privacy Protection

### v3

@@ 3.2.4 Quantization-Based Approach
The O-PQ algorithm also updates the codebooks and codes for each subvector using an online learning technique. The algorithm uses two parameters: $\alpha$, which controls the learning rate, and $\beta$, which controls the forgetting rate. The algorithm updates the codebooks and codes as follows: For each new point $x$, assign it to its nearest centroid in each subvector using PQ. For each subvector $x_{i}$, update its centroid $c_{q_{i}\left(x_{i}\right)}$ as:
@@
The advantage of O-PQ is that it can deal with changing data distributions and new data points, as it can update the codebooks and the codes in real time. O-PQ search performance depends on roughly the same factors as OPQ and also faces similar challenges as OPQ.
Scalable Nearest Neighbor. It is a technique for efficient vector similarity search at scale [guo2020accelerating, sun2023soar]. ScaNN optimizes Maximum Inner Product Search (MIPS) through search space pruning and quantization. Traditional MIPS schemes aim to minimize the average distance between each vector $x$ and its centroids $\tilde{x}$, that is, to minimize quantization distortions. The formula for typically measuring the quantization distortion is:
@@
where $N$ is the total number of vectors, $x_{i}$ is the original vector, $\tilde{x}$ is quantized centroid, and $\|\cdot\|_{2}$ denotes Euclidean norm.
@@
ScaNN takes into account the distribution characteristics of the data in different directions; ellipsoidal or other shaped regions are used instead of spherical regions around the centroids to better fit the local structure of the data. Building on this perspective, the anisotropic loss function can further enhance the adaptability of vector quantization to data anisotropy. By explicitly separating quantization errors into parallel and orthogonal components, the anisotropic loss function assigns distinct scaling parameters $h_{i}^{\|}$ and $h_{i}^{\perp}$ to these components, respectively. This allows for more fine-grained control over the quantization process, ensuring that the errors are distributed in alignment with the data’s geometric characteristics.
@@
By integrating this anisotropic loss framework, the quantization process moves beyond spherical symmetry and better accommodates ellipsoidal or irregularly shaped distributions in the data. The performance of ScaNN depends on several factors, such as the anisotropy of the data distribution, the choice of quantization methods like vector or product quantization, the size and quality of the codebooks, and the efficiency of the partitioning and scoring processes.
@@
Inverted File Product quantization. It is a widely used technique for approximate nearest neighbor (ANN) search in high-dimensional vector spaces [guo2022manu, wang2021milvus]. This algorithm is a combination of the Inverted File Indexing (IVF) and Product Quantization (PQ) algorithms. IVF_PQ first uses the IVF algorithm to divide or partition the data into clusters and uses the parameter nprobenprobe to control the number of clusters. The higher the nprobenprobe, the better the search results, but it also increases the time required. It then identifies the top-N clusters closest to the query vector and performs the search within these N clusters using the Product Quantization (PQ) algorithm.
The IVF_PQ algorithm naturally results in two different approaches when using the PQ algorithm: the first involves performing K-means clustering with the IVF algorithm, followed by applying a local PQ algorithm for dimensionality reduction within each cluster; the second also starts with the IVF algorithm to divide all data points into several clusters but applies a globally unified PQ algorithm for dimensionality reduction within each cluster.
Randomized Bit Quantization[gao2024rabitq]. It is a recently proposed quantization method that addresses a key limitation of existing approaches: the lack of theoretical guarantees on distance estimation error, which can lead to catastrophic failures on some real-world datasets. It compresses D-dimensional vectors into D-bit strings while providing a sharp theoretical bound on the approximation error, ensuring that distance estimates are provably close to true distances. Beyond its theoretical guarantees, RaBitQ achieves excellent empirical performance through efficient implementations leveraging bitwise operations or SIMD-based acceleration. Extensive experiments show that RaBitQ outperforms PQ and its variants in accuracy-efficiency trade-off by a clear margin, with empirical behavior closely aligning with theoretical analysis. This combination of rigorous guarantees and practical efficiency makes RaBitQ particularly valuable for applications requiring both reliability and high performance.
3.2.5 Summary and Trade-offs of ANNS Methods
Four mainstream approximate nearest neighbor search methods exist: hashing-based, tree-based, graph-based, and quantization-based approaches. Each offers different trade-offs in accuracy, speed, memory, and dynamic update support.
Hashing-based methods (e.g., LSH) suit static high-dimensional sparse data but often have limited recall. Tree-based methods (e.g., KD-Tree) perform well in low dimensions but degrade in high dimensions due to the curse of dimensionality. Graph-based methods (e.g., HNSW) balance accuracy and speed effectively, making them popular in production, though they require high memory and careful maintenance. Quantization-based methods (e.g., PQ) compress vectors to reduce storage, ideal for memory-constrained settings, but may sacrifice accuracy.
Choosing the right method requires balancing data dimensionality, query patterns, hardware resources, and update frequency.
4 Vector Database Comparison
@@
4.1 The Comparison of Features and Characteristics of Vector Databases
The characteristics of VDBs directly affect their performance in practical applications. Therefore, gaining a deep understanding of these databases’ features is essential for selecting the most suitable one. As shown in Table 1, we compare several popular VDBs, focusing on their differences in indexing methods, query types, distance functions, scalability, maximum dimension, and support for data management features such as replication, sharding, and partitioning.
It can be observed from Table  1 that all VDBs support NNS and ANNS. However, the implementation strategies and optimizations for these searches vary significantly across databases, depending on their underlying indexing methods and architectural designs. For example, the indexing methods and distance functions are not exactly the same across databases, but there are commonalities. For instance, all databases except Pinecone support graph-based methods, which indicates that graph-based methods are widely adopted for their ability to handle complex relationships and data structures. Additionally, the majority of databases also support three distance functions: inner product, cosine similarity, and Euclidean distance. For details on the indexing methods and distance functions supported by different databases, see Table 2 and Table 3 below.
Table 1: FEATURES OF VECTOR DATABASES
@@ 4.1 The Comparison of Features and Characteristics of Vector Databases
Table 2: Overview of Supported Distance Functions in VDBs
@@
Table 3: Overview of Supported Indexing Methods in VDBs
@@
SPARSE Inverted Index $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ N/A
SPARSE WAND $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ N/A
@@
The last column of table 1 provides statistics on the maximum vector dimensions supported by each database. It can be observed that, with the exception of ElasticCloud and Pinecone, for which no relevant information was available, most of the listed VDBs support a total vector dimension in the range of tens of thousands, with the maximum supported dimensions ranging from 16,000 to 65,535. It should be noted that QdrantCloud has default support for up to 65,535 dimensions, though this can be configured to support higher dimensions.
4.2 The Comparison of Loading Capacity and Search Performance of VDB
@@
VectorDBBench provides a comprehensive performance analysis by evaluating VDBs based on metrics such as Queries Per Second (QPS), recall rate, latency (the time required for each query from submission to system response), load duration, and maximum load count (The maximum number of vectors a database can successfully insert or store in a single loading operation). Its testing methodology employs a relative scoring mechanism to ensure fair comparisons. For QPS, the highest observed value among all tested databases serves as the reference baseline; for latency, the lowest observed value among all tested databases is used as the baseline, with an additional 10 ms adjustment to avoid distortions when latency is very low. For systems that fail or encounter timeouts in a specific test case, their scores are penalized by assigning a value proportionally worse than the lowest-performing result, using a factor of two. For example, in the case of QPS, the score is reduced to half of the minimum observed value, while for latency, it is increased to twice the maximum observed value. The formulas for calculating QPS and latency metrics for VDB $x$ are as follows:
@@
where $origin\_QPS_{x}$ and $origin\_Latency_{x}$ represent the original QPS value and original latency value, respectively, measured for database $x$ during the test. base_QPSbase\_QPS and base_Latencybase\_Latency is the reference baseline.
Table 4: VECTOR DATABASE EVALUATION TEST CASES
@@ 3 Search Performance Google C43 500K 1536 N/A IBT, R, L, MQPS
Specifically, as shown in Table 4, the VDB evaluation consists of a series of test cases designed to assess capacity, search performance, and filtering search performance. Capacity cases (Cases 1 and 2) measure the database’s ability to handle large datasets, focusing on the number of inserted vectors using SIFT and GIST datasets. Search performance cases (Case 3) evaluate index building time, recall, latency, and maximum QPS using the Google C4 dataset.
Figure 7: Performance Test Results for VDBs
@@
The VDB versions involved in the performance tests are as follows: Milvus-2c8g-hnsw-v2.2.12 (hereafter referred to as Milvus), Pinecone-p1.x1 (hereafter referred to as Pinecone), WeaviateCloud-standard (hereafter referred to as Weaviate Cloud), ZillizCloud-2cu-cap-v2023.6 (hereafter referred to as ZillizCloud), QdrantCloud-2c8g-1node (hereafter referred to as QdrantCloud), PgVector-2c8g (hereafter referred to as PgVector), and ElasticCloud-upTo2.5c8g (hereafter referred to as ElasticCloud). To ensure minimal differences in hardware performance across the tested databases, a configuration of 2 CPUs and 8GB of memory was specifically selected. For VDBs that do not meet this hardware requirement, similar configurations were chosen as closely as possible. The specific test results are shown in the Figure 7. The overall ranking in Figure 7 is calculated by averaging the rankings of each sub-test item, with the final overall ranking determined in ascending order of the average values. According to the overall ranking, Milvus ranks first with its comprehensive performance, achieving a QPS of 380, latency of 12.4 milliseconds, and outstanding performance in load capacity. In contrast, ElasticCloud ranks last, with a QPS of 11.29 and latency as high as 361 milliseconds. In terms of recall rate, all databases perform close to 1.0, indicating little difference in query accuracy. ZillizCloud demonstrates the best latency performance, at only 6 milliseconds, but its load capacity is relatively low. It is worth noting that ZillizCloud, ElasticCloud, and PgVector were assigned penalty values (P) due to timeouts in the load capacity tests, which may have been caused by network issues and should not be taken as a definitive measure of their actual performance. Overall, no single database ranked in the top three across all tests, indicating that different databases may have their own strengths. Additionally, the overall rank is simply a straightforward average of the rankings in each test item, without applying weighted averages based on specific tasks, so it should be regarded as a preliminary reference.
A deeper analysis of these performance results reveals that multiple factors influence the outcomes, including hardware configuration, deployment mode, index parameter settings, and data distribution. In this benchmark, all databases were configured with 2 CPU cores and 8 GB of memory to minimize hardware discrepancies. However, different databases exhibit varying efficiency in resource utilization. For instance, ZillizCloud demonstrated outstanding low-latency performance but faced limitations in load capacity due to network timeouts and indexing strategies; Milvus excelled in both QPS and load capacity, likely due to its efficient index implementation and parallel processing capabilities. It is important to note the typical trade-off between QPS and recall: higher query throughput often comes at the cost of lower recall. In this evaluation, Pinecone achieved a relatively low recall (0.8064) but still maintained competitive QPS, suggesting that it may prioritize speed over accuracy in certain scenarios. The timeouts observed in PgVector and ElasticCloud during load tests may be attributed to data ingestion methods, network conditions, or index construction efficiency, and should not be interpreted as definitive indicators of their performance. Therefore, when selecting a VDB, users should consider their specific application requirements—such as real-time responsiveness, data scale, and hardware constraints—and are encouraged to conduct customized evaluations in their own environments to validate performance.
Beyond the established vector databases evaluated above, recent research has introduced HARMONY, a distributed ANNS system that addresses load imbalance and high communication overhead caused by traditional partition strategies[xu2025harmony]. It uses multi-granularity partitioning to balance load and minimize communication, plus an early-stop pruning mechanism to reduce overhead. Experiments show HARMONY achieves 4.63× higher throughput than leading distributed vector databases and 58% improvement for skewed workloads, offering a promising direction for future VDB development.
5 Challenges
5.1 High-Dimensional Vector Index and Search
@@
VDBs require efficient indexing and searching of billions of vectors in hundreds or thousands of dimensions, which poses a huge computational and storage challenge. Traditional indexing methods, such as B-trees or hash tables, are not suitable for high-dimensional vectors because they suffer from dimensionality catastrophe. Therefore, VDBs need to use specialized techniques such as ANN search, hashing, quantization, or graph-based search to reduce complexity and improve the accuracy of vector similarity search.
5.2 Support for Heterogeneous Vector Data Types
@@
5.3 Distributed Parallel Processing Support
@@
5.4 Integration with Emerging Application Scenarios
@@
Currently, many emerging application scenarios remain underexplored. For instance, the incremental k-NN search [zhao2008incremental] adopted by recommendation and e-commerce platforms faces significant challenges due to the vast imbalance between the volume of vector data processed and the data displayed to users. This method cannot be effectively supported by most VDBs [pan2024survey]. Furthermore, with the latest advancements in sparse vector technology, integrating these technologies into VDBs to enable hybrid retrieval (combining keyword and vector retrieval methods) is increasingly regarded as a best practice. Such hybrid systems must manage large datasets while enhancing computational efficiency and maintaining retrieval quality.
5.5 Data Security and Privacy Protection